# Taming the Long-Tail: Efficient Reasoning RL Training with Adaptive Drafter

### v1

@@ 5.2. Auto-Tune Algorithm
Bucketed -Greedy MAB. Multi-Armed Bandit (MAB) algorithms are a classical framework for online decision-making, excelling at maximizing cumulative gains when selecting actions (“arms”) with uncertain rewards (MABResourceAllocation; MAB). Each ”arm” corresponds to a specific speculative decoding configuration tuple: (Draft_Depth, topK, Tokens_to_Verify), and the ”reward” reflects the efficiency within the generation step (i.e., ).
@@ 6.1. Experimental Setup
Implementation. We implement TLT on top of VeRL (HybridFlow) and utilizes Ray (Ray) for distributed execution. The Worker Coordinator is realized with a centralized controller that employs ZeroMQ (ZeroMQ) messaging. The Adaptive Drafter’s implementation is based on training pipelines from EAGLE and trained with FSDP2 (FSDP2). For asynchronous checkpointing, we leverage PyTorch DCP (DCP). To further improve generation efficiency, we employ SGLang (SGLang) as the rollout backend, with implementing adaptive enabling of SD, drafter weight update and the BEG-MAB tuner following (SGLangMAB). We additionally use (LookaheadKDD) as our model-free drafter.
Testbed. We conduct experiments on a cluster of 8 NVIDIA DGX H100 (dgx-h100) servers, totaling 64 GPUs. Each server is equipped with 8 NVIDIA H100 GPUs, 2 Intel Xeon Platinum 8480C (112 cores) CPUs and 2TB memory. GPUs are interconnected to each other by 900GB/s NVLink, and inter-node communication is achieved via NVIDIA Mellanox 400Gb/s InfiniBand. To evaluate system’s performance across heterogeneous hardware, we also conduct experiments on a different cluster equipped with NVIDIA DGX A100 servers. Unless otherwise specified, experiments are conducted on H100 GPUs by default.
Models. Following the training procedure introduced in DeepSeek-R1 (DeepSeekR1), the reasoning RL process can begin either from a fine-tuned model or directly from a base model without additional fine-tuning. To demonstrate that TLT robustly accommodates different model architectures, scales, and training paradigms, we evaluate the following models:
@@
Qwen2.5-32B (Qwen-32B): a larger base model without any additional fine-tuning, as in (DAPO);
@@
Datasets. For the primary RL training dataset, we use a subset from Eurus-2-RL (Prime), a comprehensive dataset curated for reasoning RL. This dataset contains mathematics and coding problems equipped with verifiers (LaTeX answers and test cases, respectively). The coding problems originate from competitive programming sources (e.g., CodeContests (CodeContests), Codeforces (codeforces)), while the math problems range from high school to olympiad-level difficulty (numina_math). For drafter training, we use a subset from OpenThoughts2-1M (OpenThoughts2-1M) as the warm-up dataset.
RL Settings. We primarily follow the reasoning RL algorithm GRPO proposed by DeepSeek (DeepSeekR1; DeepSeekMath). Specifically, for each prompt, we set the temperature=0.9, with a maximum generation length of 32K tokens, consistent with common RL training settings. We apply TP= 2, 4, or 8 depending on model scale for the rollout engine. Additionally, all experiments utilize mixed-precision training with the Adam optimizer (Adam), where BF16 precision is employed during both model rollout and inference stages.
@@
Open-R1 (openr1): a popular RL training framework built upon TRL(TRL), integrating vLLM (vLLM) for rollouts and DeepSpeed (ZeRO) for training. Currently, it only supports separate model placement, requiring serving and training to be performed on distinct nodes.
VeRL (HybridFlow): the state-of-the-art open-source RL training framework by ByteDance. It enables colocation of models on shared devices by utilizing GPU time-sharing.
@@
Metrics. Following (HybridFlow), we use token throughput (tokens/sec) for end-to-end evaluation. This metric is calculated by dividing the total number of tokens in both prompts and responses within a global batch by the duration of a single RL step. We average results over three training steps following an initial warm-up step and use a trained draft model for evaluation.
@@ 6.2. End-to-End Evaluation
We evaluate TLT against popular RL training frameworks, including Open-R1 (openr1) and VeRL (HybridFlow), across multiple GPU platforms (NVIDIA H100 and A100) and various LLM scales. Figure 11 demonstrates that TLT consistently outperforms VeRL. Open-R1 consistently underperforms, primarily because it struggles to fully utilize all available GPUs concurrently and suffers from a tight coupling between its rollout and training batch size. Similar substantial gains are observed on the A100 platform, highlighting its effectiveness across different hardware generations. We also present results for TLT-Base, which utilizes a model-free drafter and yields good benefits, demonstrating that the TLT framework provides continuous efficiency gains throughout the entire training process. These end-to-end performance gains stem directly from TLT’s ability to mitigate the long-tail rollout bottleneck of reasoning RL training.
@@ 6.3. Evaluation of Adaptive SD
In this section, we provide in-depth analysis on the effectiveness of our design for adaptive speculative decoding.
@@ 6.4. Evaluation of Spot Trainer
System Overhead Analysis. We identify three primary sources of overhead in TLT: (1) Stage-transition overhead, introduced by additional the drafter-weight update and optimizer offloading compared to VeRL (HybridFlow). This accounts for less than 1% of the RL-step duration due to the much smaller model size. (2) SD switch overhead, incurred when switching from normal decoding to SD. Because the draft model is initially inactive for large batch sizes, we perform a reprefill step to activate SD, which completes within 3 seconds. (3) Drafter-training coordination overhead, caused by occasional contention during spot-trainer execution. However, we find that it is unnecessary to train the drafter every RL step, performing once every 10 steps is sufficient to maintain accuracy. Overall, the performance gains of TLT far outweigh this overhead, yielding end-to-end speedups.
@@ 7. Discussion
Can We Break RL Synchronization? TLT aims to improve RL training efficiency without modifying on-policy requirements of the underlying algorithm. One may wonder whether it is possible to further accelerate training by asynchronously updating the model with partial rollouts during long decoding phases. While appealing, this approach risks altering the underlying RL algorithm and potentially harming convergence and model performance. Asynchronous RL typically mixes multiple policy versions within a single rollout, potentially biasing gradient estimation and shifting the learning dynamics (e.g., drifting from on-policy to off-policy). Additional algorithmic modifications are required to safely accommodate such paradigm (AReal). A more balanced approach is to allow limited asynchrony that preserves algorithmic correctness while improving hardware utilization. Even under such settings, the heavy-tail rollout problem remains, and our speculative decoding approach can be combined with limited asynchronous RL to further accelerate training without introducing additional loss or policy staleness, leaving this as an intriguing direction for future work.
Can TLT be applied to other RL algorithms? GRPO represents a widely used algorithmic structure for reasoning RL. Alternatives such as RLOO (RLOO), DAPO (DAPO), REINFORCE (REINFORCE), and REINFORCE++ (REINFORCEpp) generally share the same overall training workflow, differing mainly in their reward formulations and use of KL-regularization. These algorithmic variations remain fully compatible with our adaptive drafter and spot-training design, suggesting that TLT can readily accelerate a broad class of RL methods.
@@ 8. Related Work
RL Systems for LLMs. Several systems aim to optimize the training pipeline of Reinforcement Learning from Human Feedback (RLHF) (InstructGPT). Frameworks like DeepSpeed-Chat (DeepSpeedChat), GEAR (GEAR), VeRL (HybridFlow), FlexRLHF (FlexRLHF) and NeMo-Aligner (NeMoAligner) provide scalable end-to-end solutions, supporting various alignment algorithms and parallelism techniques. Others focus on specific optimizations within the pipeline. RLHFuse (RLHFuse) employs fine-grained scheduling and stage fusion to improve GPU utilization. ReaLHF (ReaLHF) introduces dynamic parameter reallocation to flexibly adapt 3D parallelism strategies across RLHF stages. Their optimization efforts primarily target the efficient management of multiple models inherent in the RLHF process. However, they neglect a critical performance bottleneck in the rollout phase. This oversight renders existing systems inefficient for reasoning RL tasks. Our work addresses this gap by concentrating specifically on optimizing the rollout bottleneck, utilizing adaptive drafter which is orthogonal to prior systems. Recent efforts (AReal; StreamRL) attempt to alleviate the synchronization constraint in RL by allowing stale responses for model updates. While such asynchronous strategies accelerate training, they risk degrading model quality. In contrast, our approach preserves the original RL algorithm, ensuring lossless training performance while achieving significant efficiency gains.
@@
Speculative Decoding. Speculative decoding (SpeculativeDecoding; chen2023accelerating; AdaServe) accelerates LLM inference by verifying candidate tokens in parallel with the target model, effectively increasing the number of tokens generated per step. Enhancements to speculative decoding encompass diverse strategies. Researchers have explored retrieval-based methods (REST; saxena2023PLD; lookahead), which avoid the need to train an extra draft model. Others modify the target model directly: Medusa (Medusa) integrates multiple prediction heads, an approach refined by Hydra (Hydra) which models correlations between these predictions. Furthermore, SpecInfer (SpecInfer) innovatively employs tree-based speculative decoding to broaden the candidate search space. OSD (OSD) leverages knowledge distillation for the drafter model so it remains aligned with distribution shifts in online serving. Additionally, EAGLE (Eagle1; Eagle2; Eagle3) trains separate, compact auto-regressive models as dedicated draft models. Further refining this approach, HASS (HASS) improves draft model accuracy by simulating multi-step predictions during training, mitigating the training-inference inconsistency issue. Recent work also adapts speculative idea for efficient reasoning LLM inference. For example, RSD (RSD) employs a reward model to guide the speculation, and SpecReason (SpecReason) uses a lightweight model to speculatively execute intermediate reasoning steps. Unlike these approaches designed for inference with static models, we focus on the dynamic RL training scenario where the target model is continuously updated.
@@ 9. Conclusion
To conclude, TLT alleviates the critical long-tail bottleneck in reasoning RL training through an adaptive drafter. The adaptivity operates on two folds: during training, it adapts to target model updates; during inference, it adapts to varying batchsizes. As a result, TLT achieves substantial throughput improvements while preserving model quality.
@@ Acknowledgements
We thank MIT-IBM Watson AI Lab, MIT AI Hardware Program, MIT Amazon Science Hub, Hyundai Motor Company, and National Science Foundation for supporting this work.

### v2

@@ 5.2. Auto-Tune Algorithm
Bucketed -Greedy MAB. Multi-Armed Bandit (MAB) algorithms are a classical framework for online decision-making, excelling at maximizing cumulative gains when selecting actions (“arms”) with uncertain rewards (Thananjeyan et al., 2021; Slivkins and others, 2019). Each ”arm” corresponds to a specific speculative decoding configuration tuple: (Draft_Depth, topK, Tokens_to_Verify), and the ”reward” reflects the efficiency within the generation step (i.e., ).
@@ 6.1. Experimental Setup
Implementation. We implement TLT on top of VeRL (Sheng et al., 2025) and utilizes Ray (Moritz et al., 2018) for distributed execution. The Worker Coordinator is realized with a centralized controller that employs ZeroMQ (Hintjens, 2013) messaging. The Adaptive Drafter’s implementation is based on training pipelines from EAGLE and trained with FSDP2 (PyTorch Team, 2025a). For asynchronous checkpointing, we leverage PyTorch DCP (PyTorch Team, 2025b). To further improve generation efficiency, we employ SGLang (Zheng et al., 2024) as the rollout backend, with implementing adaptive enabling of SD, drafter weight update and the BEG-MAB tuner following (Su, 2025). We additionally use (Zhao et al., 2024) as our model-free drafter.
Testbed. We conduct experiments on a cluster of 8 NVIDIA DGX H100 (NVIDIA Corporation, 2025) servers, totaling 64 GPUs. Each server is equipped with 8 NVIDIA H100 GPUs, 2 Intel Xeon Platinum 8480C (112 cores) CPUs and 2TB memory. GPUs are interconnected to each other by 900GB/s NVLink, and inter-node communication is achieved via NVIDIA Mellanox 400Gb/s InfiniBand. To evaluate system’s performance across heterogeneous hardware, we also conduct experiments on a different cluster equipped with NVIDIA DGX A100 servers. Unless otherwise specified, experiments are conducted on H100 GPUs by default.
Models. Following the training procedure introduced in DeepSeek-R1 (DeepSeek-AI et al., 2025), the reasoning RL process can begin either from a fine-tuned model or directly from a base model without additional fine-tuning. To demonstrate that TLT robustly accommodates different model architectures, scales, and training paradigms, we evaluate the following models:
@@
Qwen2.5-32B (Qwen-32B): a larger base model without any additional fine-tuning, as in (Yu et al., 2025);
@@
Datasets. For the primary RL training dataset, we use a subset from Eurus-2-RL (Cui et al., 2025), a comprehensive dataset curated for reasoning RL. This dataset contains mathematics and coding problems equipped with verifiers (LaTeX answers and test cases, respectively). The coding problems originate from competitive programming sources (e.g., CodeContests (Lewkowycz et al., 2022), Codeforces (Codeforces, 2025)), while the math problems range from high school to olympiad-level difficulty (LI et al., 2025). For drafter training, we use a subset from OpenThoughts2-1M (OpenThoughts, 2025) as the warm-up dataset.
RL Settings. We primarily follow the reasoning RL algorithm GRPO proposed by DeepSeek (DeepSeek-AI et al., 2025; Shao et al., 2024). Specifically, for each prompt, we set the temperature=0.9, with a maximum generation length of 32K tokens, consistent with common RL training settings. We apply TP= 2, 4, or 8 depending on model scale for the rollout engine. Additionally, all experiments utilize mixed-precision training with the Adam optimizer (Kingma and Ba, 2015), where BF16 precision is employed during both model rollout and inference stages.
@@
Open-R1 (Face, 2025): a popular RL training framework built upon TRL(von Werra et al., 2025), integrating vLLM (Kwon et al., 2023) for rollouts and DeepSpeed (Rajbhandari et al., 2020) for training. Currently, it only supports separate model placement, requiring serving and training to be performed on distinct nodes.
VeRL (Sheng et al., 2025): the state-of-the-art open-source RL training framework by ByteDance. It enables colocation of models on shared devices by utilizing GPU time-sharing.
@@
Metrics. Following (Sheng et al., 2025), we use token throughput (tokens/sec) for end-to-end evaluation. This metric is calculated by dividing the total number of tokens in both prompts and responses within a global batch by the duration of a single RL step. We average results over three training steps following an initial warm-up step and use a trained draft model for evaluation.
@@ 6.2. End-to-End Evaluation
We evaluate TLT against popular RL training frameworks, including Open-R1 (Face, 2025) and VeRL (Sheng et al., 2025), across multiple GPU platforms (NVIDIA H100 and A100) and various LLM scales. Figure 11 demonstrates that TLT consistently outperforms VeRL. Open-R1 consistently underperforms, primarily because it struggles to fully utilize all available GPUs concurrently and suffers from a tight coupling between its rollout and training batch size. Similar substantial gains are observed on the A100 platform, highlighting its effectiveness across different hardware generations. We also present results for TLT-Base, which utilizes a model-free drafter and yields good benefits, demonstrating that the TLT framework provides continuous efficiency gains throughout the entire training process. These end-to-end performance gains stem directly from TLT’s ability to mitigate the long-tail rollout bottleneck of reasoning RL training.
@@ 6.3. Evaluation of Adaptive SD
GPU Diversity and Scalability. We evaluate TLT across diverse GPU architectures and cluster scales. As in Table 2, TLT delivers consistent speedups on both data-center and consumer GPUs. Furthermore, end-to-end training results (Table 3) demonstrate that TLT’s acceleration becomes more pronounced as model size and cluster size increase, reflecting favorable scalability for large-scale reasoning RL training.
| GPU Type | w/ SD | w/o SD | Speedup |
| B200 | 605.05 | 259.71 | 2.33 |
| H100 | 430.24 | 164.65 | 2.61 |
| A100 | 259.01 | 92.83 | 2.79 |
| RTX 5090 | 293.84 | 100.89 | 2.91 |
| RTX 4090 | 187.44 | 65.28 | 2.87 |
| RTX 3090 | 166.41 | 51.75 | 3.22 |
| Nodes (8 GPUs/Node) | 1 Node | 2 Nodes | 4 Nodes | 8 Nodes |
| Qwen2.5-7B | 1.21 | 1.45 | 1.62 | 1.76 |
| Qwen2.5-32B | OOM | OOM | 1.83 | 2.12 |
@@ 6.4. Evaluation of Spot Trainer
System Overhead Analysis. We identify three primary sources of overhead in TLT: (1) Stage-transition overhead, introduced by the additional drafter-weight update and optimizer offloading compared to VeRL (Sheng et al., 2025). This accounts for less than 1% of the RL-step duration due to the much smaller model size. (2) SD switch overhead, incurred when switching from normal decoding to SD. Because the draft model is initially inactive for large batch sizes, we perform a reprefill step to activate SD, which completes within 3 seconds. (3) Drafter-training coordination overhead, caused by occasional contention during spot-trainer execution. However, we find that it is unnecessary to train the drafter every RL step, performing once every 10 steps is sufficient to maintain accuracy. Overall, the performance gains of TLT far outweigh this overhead, yielding end-to-end speedups.
6.5. Analysis on Drafter Architectures
To demonstrate the generality of the TLT framework, we integrate and evaluate several state-of-the-art speculative decoding methods in Table 7. While newer SD methods like HASS (Zhang et al., 2025) and Eagle-3 (Li et al., 2025a) offer slightly higher accept length and throughput, their ”Training-Time Test” technique, which requires multiple forward passes during the training process, significantly increases the training overhead per step and the difficulty of model convergence. Given the limited time budget in the rollout bubble, we chose Eagle (Li et al., 2024a) for its comparable performance and simplicity.
6.6. Discuss Online SD (OSD)
OSD (Liu et al., 2024) leverages general knowledge distillation (GKD) for the drafter model so it remains aligned with distribution shifts in online serving. In contrast, TLT focuses on RL training, and the target model’s weights are continuously evolving. While KD is already incorporated as a loss term in Eagle training, OSD further explores reverse KD. We investigate the impact of different training paradigms in Table 8. The results demonstrate that OSD-style training more effectively improves draft model quality than supervised fine-tuning (SFT) for a general small model (e.g., Qwen-0.5B), as well as the default Eagle training method used in TLT.
| Qwen2.5-0.5B | 5.95 | 363.87 | 6.68 | 408.51 | 6.89 | 421.35 | ||||||||||||||
| Eagle | 1.57 | 130.34 | 6.53 | 542.12 | 6.77 | 562.04 |
@@ 7. Discussion
Can We Break RL Synchronization? TLT aims to improve RL training efficiency without modifying on-policy requirements of the underlying algorithm. One may wonder whether it is possible to further accelerate training by asynchronously updating the model with partial rollouts during long decoding phases. While appealing, this approach risks altering the underlying RL algorithm and potentially harming convergence and model performance. Asynchronous RL with or continuous batching (Fu et al., 2025) typically mixes multiple policy versions within a single rollout, potentially biasing gradient estimation and shifting the learning dynamics (e.g., drifting from on-policy to off-policy). Additional algorithmic modifications are required to safely accommodate such paradigm. A more balanced approach is to allow limited asynchrony that preserves algorithmic correctness while improving hardware utilization. Even under such settings, the heavy-tail rollout problem remains, and our speculative decoding approach can be combined with limited asynchronous RL to further accelerate training without introducing additional loss or policy staleness.
Can TLT be applied to other RL algorithms? GRPO represents a widely used algorithmic structure for reasoning RL. Alternatives such as RLOO (Ahmadian et al., 2024), DAPO (Yu et al., 2025), REINFORCE (Williams, 1992), and REINFORCE++ (Hu et al., 2025) generally share the same overall training workflow, differing mainly in their reward formulations and use of KL-regularization. These algorithmic variations remain fully compatible with our adaptive drafter and spot-training design, suggesting that TLT can readily accelerate a broad class of RL methods.
@@ 8. Related Work
RL Systems for LLMs. Several systems aim to optimize the training pipeline of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022). Frameworks like (Yao et al., 2023; Wang et al., 2023; Sheng et al., 2025; Xiao et al., 2024b; Shen et al., 2024) provide scalable end-to-end solutions, supporting various alignment algorithms and parallelism techniques. RLHFuse (Zhong et al., 2025b) employs fine-grained scheduling and stage fusion to improve GPU utilization. ReaLHF (Mei et al., 2024) introduces dynamic parameter reallocation to flexibly adapt 3D parallelism strategies across RLHF stages. Their optimization efforts primarily target the efficient management of multiple models inherent in the RLHF process. However, they neglect a critical performance bottleneck in the rollout phase. This oversight renders existing systems inefficient for reasoning RL tasks. Our work addresses this gap by concentrating specifically on optimizing the rollout bottleneck, utilizing an adaptive drafter that is orthogonal to prior systems. Recent efforts (Fu et al., 2025; Zhong et al., 2025a; Zhou et al., 2025) attempt to alleviate the synchronization constraint in RL by allowing stale responses for model updates or terminating rollout prematurely. RollPacker (Gao et al., 2025a) introduces tail batching; however, even for identical queries, reasoning trajectories can vary substantially in length, limiting its effectiveness. While these strategies accelerate training, they risk degrading model quality. In contrast, our approach preserves the original RL algorithm, ensuring lossless training performance while achieving significant efficiency gains.
@@
Speculative Decoding. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Li et al., 2025b) accelerates LLM inference by verifying candidate tokens in parallel with the target model, effectively increasing the number of tokens generated per step. Enhancements to speculative decoding encompass diverse strategies. Researchers have explored retrieval-based methods (He et al., 2024; Saxena, 2023; Fu et al., 2024), which avoid the need to train an extra draft model. Others modify the target model directly: The Medusa series (Cai et al., 2024) integrates multiple prediction heads, an approach refined by Hydra (Ankner et al., 2024) which models correlations between these predictions. Furthermore, SpecInfer (Miao et al., 2024) innovatively employs tree-based speculative decoding to broaden the candidate search space. OSD (Liu et al., 2024) leverages knowledge distillation for the drafter model so it remains aligned with distribution shifts in online serving. Additionally, EAGLE (Li et al., 2024b, a, 2025a) trains separate, compact auto-regressive models as dedicated draft models. Further refining this approach, HASS (Zhang et al., 2025) improves draft model accuracy by simulating multi-step predictions during training, mitigating the training-inference inconsistency issue. Recent work also adapts speculative ideas for efficient reasoning LLM inference. For example, RSD (Liao et al., 2025) employs a reward model to guide the speculation, and SpecReason (Pan et al., 2025) uses a lightweight model to speculatively execute intermediate reasoning steps. Unlike these approaches designed for inference with static models, we focus on the dynamic training scenario where the target model is continuously updated.
@@ 9. Conclusion
To conclude, TLT alleviates the long-tail bottleneck in reasoning RL training via an adaptive drafter, delivering substantial throughput gains without compromising model quality.
@@ Acknowledgements
We sincerely thank our shepherd, Prof. Jiarong Xing, for his guidance throughout the paper revision process. We also thank the anonymous reviewers of ASPLOS and SOSP for their insightful and valuable feedback. We thank MIT-IBM Watson AI Lab, MIT AI Hardware Program, MIT Amazon Science Hub, Hyundai Motor Company, and National Science Foundation for supporting this work.